OpenStudy (anonymous):

(5,7), (-6,-6) distance formula

OpenStudy (anonymous):

((-6-7)^2+(-6-5)^2)^1/2=4.13

OpenStudy (anonymous):

\[distance=\sqrt{\Delta x^2+\Delta y^2}\] So,\[distance =\sqrt{(-6-5)^2+(-6-7)^2}\]We have,\[distance = \sqrt{121+169}=\sqrt{290}\]

OpenStudy (anonymous):

\[\sqrt{(5+6)^2+(7+6)^2}\]